OpenStudy (anonymous):

one kilogram is about 2.205 pounds. When a 55 gallon aquarium is filled, it weights 625 lb. to the nearest tenth, how many kilograms os this

OpenStudy (anonymous):

If you know 1 Kg = 2.205 lbs. This is represented as \[\frac{ 1 Kg }{ 2.205 lbs }\] So to get Kg from a known lbs you can write this as: \[\frac{ 1 kg }{ 2.205 lbs } *\frac{ 625lbs }{ 1 kg } ==\frac{ 625 lbs }{ 2.205 lbs } = x \]